OpenStudy (anonymous):

A quadratic equation is shown below: x^2 - 14x + 41 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)^2 = q, where p and q are integers?

OpenStudy (anonymous):

Add 8 to both sides of the equation Add 9 to both sides of the equation Subtract 8 from both sides of the equation Subtract 9 from both sides of the equation

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

Hint:- you need to make the left side a perfect square so the last term must be a perfect square

OpenStudy (welshfella):

- also when the trinomial is expanded the second term must be -14x

OpenStudy (anonymous):

add 8 to both sides?

OpenStudy (welshfella):

right - to give 49 as the last term on LHS which will also give the -14x

OpenStudy (anonymous):

thanks!

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

What are the exact solutions of x^2 - 3x - 1 = 0?

OpenStudy (welshfella):

this is an exercise in completing the square first divide the coefficient of x ( that us the -3) by 2 and write x^2 - 3x - 1 = ( x - 1.5)^2 -2.25 - 1 = 0 the -2.25 comes from when you square -1.5 it becomes 2.25 so you have to subtract 2.25 to square things up

OpenStudy (welshfella):

so (x - 1.5)^2 = 3.25

OpenStudy (anonymous):

my answer was the first one

OpenStudy (welshfella):

Oh they want the answer as a fraction so i'll change it to fractions.

OpenStudy (welshfella):

lets check it....

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

no i dont get that lets see x - 3/2 = +/- sqrt13/2 = 3/2 + sqrt13/2 or 3/2 - sqrt13/2 = 3 +/- sqrt13 ---------- 2

OpenStudy (anonymous):

so C?

OpenStudy (welshfella):

yes